OpenStudy (anonymous):

HELP! Find the point-slope equation of the line using the point (1, 2) and slope of 3/2 . Use the slash key (/) to indicate a fraction.

OpenStudy (anonymous):

K ! Use this : Y2-Y=m(X2-X) Got it ? x and y are in the equation and : y1=2 x1=2

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

@E.ali 3x-y-4=0 am I right?

OpenStudy (anonymous):

Got it ?

OpenStudy (anonymous):

For this ?

OpenStudy (anonymous):

yes! that's the answer I got. I'm asking am I correct

OpenStudy (anonymous):

We have : 2-y=3/2(2-x) 2-y=3-3/2x -y=3/2x-3+2 -y=3/2x-1 y=-3/2x+1 Got it ?

OpenStudy (anonymous):

oooh! okay thanks I understand now

OpenStudy (anonymous):

You re welcome ! Any question ?!

OpenStudy (anonymous):

Nope! that'll be all

OpenStudy (anonymous):

K !